Find the volume of the solid with cross-sectional area 2/√(1−4x^2) lying between x=0 and x=12
1
Expert's answer
2020-05-24T17:13:10-0400

Let "S" be a solid and suppose that the area of the cross section in the plane perpendicular to the

 "x-"axis is "A(x)" for "a\leq x\leq b."


"V=\displaystyle\int_{a}^bA(x)dx"

"V=\displaystyle\int_{0}^{1/2}{2\over \sqrt{1-4x^2}}dx=\displaystyle\int_{0}^{1/2}{1\over \sqrt{{1 \over 4}-x^2}}dx"

Trigonometric substitution


"x=\dfrac{1}{2}\sin t, -\dfrac{\pi}{2}\leq t\leq \dfrac{\pi}{2}"


"dx=\dfrac{1}{2}\cos t\ dt"


"\sqrt{{1 \over 4}-x^2}=\dfrac{1}{2}\cos t"


"\int{1\over \sqrt{{1 \over 4}-x^2}}dx=\int{\dfrac{1}{2}\cos t\over \dfrac{1}{2}\cos t}dt=t+C=\arcsin(2x)+C"

"V=\displaystyle\int_{0}^{1/2}{2\over \sqrt{1-4x^2}}dx=\displaystyle\int_{0}^{1/2}{1\over \sqrt{{1 \over 4}-x^2}}dx="

"=[\arcsin(2x)]\begin{matrix}
 1/2 \\
 0
\end{matrix}={\pi \over 2}-0={\pi \over 2} (units^3)"

"V={\pi \over 2}\ cubic\ units"